OpenStudy (anonymous):

How to simplify the following expression : A'BCD + AB'CD' + AB'CD + ABC'D + ABCD' + ABCD ? It should get AC + BCD + ABD using Kmap but using boolean algebra i am stuck no matter how i try .

hartnn (hartnn):

u know what is A+A' = ?

OpenStudy (anonymous):

yes , it is 1

hartnn (hartnn):

so A'BCD+ABCD = ? first term + last term

OpenStudy (anonymous):

A'BCD + ABCD = BCD ( A' + A ) = BCD

hartnn (hartnn):

correct, A'BCD + AB'CD' + AB'CD + ABC'D + ABCD' + ABCD i would also add ABCD (i can because ABCD+ABCD=ABCD , right ? ) A'BCD + AB'CD' + AB'CD + ABC'D + ABCD' + ABCD + ABCD = BCD + AB'CD' + AB'CD + ABC'D + ABCD' + ABCD got this ? now ABCD+ABC'D = ?

OpenStudy (anonymous):

Yes , I got the first one ABD ( C + C' ) = ABD

hartnn (hartnn):

so now only , AC remains...let me see.....

hartnn (hartnn):

i think we need one more ABCD A'BCD + AB'CD' + AB'CD + ABC'D + ABCD' + ABCD + ABCD+ABCD = BCD + ABD + AC (B'D+BD' + BD+B'D')

hartnn (hartnn):

so now we just need to prove (B'D+BD' + BD+B'D')=1

hartnn (hartnn):

can u prove (B'D+BD' + BD+B'D')=1 ?

hartnn (hartnn):

and did u get this : A'BCD + AB'CD' + AB'CD + ABC'D + ABCD' + ABCD + ABCD+ABCD = BCD + ABD + AC (B'D+BD' + BD+B'D') ?

OpenStudy (anonymous):

i am trying to prove it ,,, n yes i got what u did so far ..

OpenStudy (anonymous):

yes it is 1

hartnn (hartnn):

(B'D+BD' + BD+B'D')=1 proving that is pretty simple, ask if u don't get it...

hartnn (hartnn):

hint : take 1st and 3rd term..... 2nd and 4th term..

OpenStudy (anonymous):

no no i can do it ... D ( B' + B ) + D' ( B' + B ) then D + D' = 1

hartnn (hartnn):

good !

hartnn (hartnn):

so clear with all steps ?

hartnn (hartnn):

the trick here was to add ABCD twice and use it..

OpenStudy (anonymous):

Yeah I was stuck becz i didn't think of adding ABCD twice and only tried to simplify with the existing ones.. Thank you soooooo much :) You are way better than my teacher ,,, by explaining easily step by step I understood really well :) Thank you so much :) You are really a lifesaver :)

hartnn (hartnn):

welcome ^_^ glad i could help :)